OpenStudy (anonymous):

Solve the differential dy/dx=(1+x)/xy for x>0 and y(1)=-4. Thanks!

myininaya (myininaya):

This can be done be separating the variables

myininaya (myininaya):

Write as g(y) dy=f(x) dx and then integrate both sides

OpenStudy (anonymous):

I have tried but failed.Can youplz howme th steps

OpenStudy (anonymous):

I appeciate it

myininaya (myininaya):

Well do you have \[\frac{dy}{dx}=\frac{1+x}{xy} ?\]

OpenStudy (anonymous):

yes

myininaya (myininaya):

Ok did you multiply dx on both sides?

OpenStudy (anonymous):

I got (x+1)/x dx=y dy

OpenStudy (anonymous):

therefore 1+ (1/x) dx=y dy

OpenStudy (anonymous):

then I integrated both sides

OpenStudy (anonymous):

I got x+ lnx=(y^2)/2

myininaya (myininaya):

(1+1/x) dx=y dy yes and yes don't forget you need +C on one of those sides

OpenStudy (anonymous):

yes

myininaya (myininaya):

so great now use the initial condition y(1)=-4 to find C

OpenStudy (anonymous):

can you me the rest

OpenStudy (anonymous):

show*

myininaya (myininaya):

You have this : \[x+\ln(x)+C=\frac{y^2}{2}\] Enter in 1 for x Enter in -4 for y Solve for C

OpenStudy (anonymous):

-_> c=7

myininaya (myininaya):

\[1+\ln(1)+C=\frac{(-4)^2}{2}\] I just replaced x with 1 and y with -4 since the initial condition says y(1)=-4 \[1+0+C=\frac{16}{2}\] \[1+C=8\] Yes C=7 very good

OpenStudy (anonymous):

I got y= -+ radical (2x+14+2lnx)

OpenStudy (anonymous):

is tht right?

myininaya (myininaya):

Yep gj! :) You multiplied 2 on both sides giving you \[2x+2\ln(x)+2C=y^2\] And we found C to be 7 So we have \[y^2=2x+2\ln(x)+14\] yes you are right :)

OpenStudy (anonymous):

Thank you so much!!!